OpenStudy (anonymous):

If q is positive and increasing, for what value of q is the rate of increase of q3 twelve times that of q?

OpenStudy (anonymous):

I am in calculus and was asked this question -- but I'm totally stumped just trying to figure out what it means. Does anyone else get it? (Answers are 2, 3, 4.5, 12, and 36) I think maybe it's 3, but I'm just not sure...

OpenStudy (anonymous):

I'm sorry, i don't get the question. What's q3?

OpenStudy (anonymous):

Is it \[q^3\]

OpenStudy (anonymous):

Oh my bad, formatting thing. Yep it's q^3

OpenStudy (anonymous):

\[(q^3)/dq = (3q^2)\]\[(q)dq = (1)\] 3q^2=12 q=2

OpenStudy (anonymous):

So then q=2 would make the rate of increase of q3 twelve times that of q?

OpenStudy (anonymous):

*q^3

OpenStudy (anonymous):

Yep That's what I get

OpenStudy (anonymous):

Thanks!